An engine is able to provide 1000 N of propulsion (applied ) force , causing a vehicle to experience an acceleration of 4.50 m/s ^ 2 If the vehicle's mass doubles and the propulsion force remains constant , what is the new acceleration ?
1
Expert's answer
2020-10-29T06:57:21-0400
"a'=\frac{F}{m'}\\a'=\frac{F}{2m}=0.5\frac{F}{m}\\a'=0.5a=0.5(4.5)=2.25\frac{m}{s^2}"